A 50hz sinusoidal voltage V=141sin Wt is supplied to a series R-L ckt having R=5ohm ,L=0.015 calculate?


1
Expert's answer
2021-03-08T07:31:59-0500
"X_L=2\pi(50)(0.015)=4.7\Omega\\Z=\sqrt{4.7^2+5^2}=6.9\Omega"